A.C. Neilsen reported that children between the ages of 2 and 5 watch an average of 25 hours of television per week. Assume the variable is normally distributed and the standard deviation is 3 hours. If 20 children between the ages of 2 and 5 are randomly selected, find the probability that the mean of the number of hours they watch television will be greater than 26.3 hours.   

Group of answer choices

14.7%

.19%

97.38%

13.1%

2.62%

Flag this Question

Question 62 1 pts

LC50 for a chemical is the lethal concentration for 50% of animals or humans. For example, if water contains a chemical at the LC50 concentration, 50% of the fish population in the water would die. Environment Canada collects data on LC50 concentrations. A 95 % confidence interval for the true mean LC50 with a random sample of 61 is (15.8, 18.4). Suppose the chemical concentration is measured in a local lake and found to be 19. This means that

Group of answer choices

at least 95% of the fish will die if nothing is done.

it is very likely that 50% of the fish will die if nothing is done.

the proability that all of the fish will dies is less than 50%.

the probability that 50% of the fish will die is more than 95%.

the probability that 95% of the fish will die is more than 5

PLEASE HURRY UP
Answer #1

#1.
Here, μ = 25, σ = 0.6708 and x = 26.3. We need to compute P(X >= 26.3). The corresponding z-value is calculated using Central Limit Theorem

z = (x - μ)/σ
z = (26.3 - 25)/0.6708 = 1.94

Therefore,
P(X >= 26.3) = P(z <= (26.3 - 25)/0.6708)
= P(z >= 1.94)
= 1 - 0.9738
= 0.0262

2.62%

2)
it is very likely that 50% of the fish will die if nothing is done.